OpenStudy (anonymous):

help plz!!!! you just got a credit card and made a prchase for $2500 the card offers a 0% apr for the first 90 days and a apr of 17.99 after wards compounded daily you dont pay off the 2500 for 1 year how much money in interest did the 0% apr for the first 90 days will save you?

OpenStudy (anonymous):

Money owed with 90 day grace period: \[$2500\times(1 + 0.1799)^{365-90}=$1.43\times10^{23}\] Money owed without grace period:\[$2500\times(1 + 0.1799)^{365}=$4.18\times10^{29}\] The 90 day grace period saved you:\[$4.18\times10^{29}-$1.43\times10^{23}=$4.18\times10^{29}\] More dollars than have ever existed, in other words. This is why interest compounded daily is absurd.